OpenStudy (anonymous):

simplify ^4root 27 (times) ^4root 3

OpenStudy (watchmath):

3

OpenStudy (anonymous):

\[\sqrt[4]{27}*\sqrt[4]{3}=\sqrt[4]{81}=\sqrt[4]{3^4}=3\]

OpenStudy (anonymous):

\[\sqrt[4]{27} * \sqrt[4]{3}\] \[\sqrt[4]{3^3 * 3}\] \[\sqrt[4]{3^4}\] \[3\]